OpenStudy (anonymous):

The base of a triangle is 5/8 feet. The height is 1/3 feet. How much greater is the base of the triangle than the height?

OpenStudy (anonymous):

5/8 = 15/24 1/3 = 8/24 7/24 is your answer.

OpenStudy (anonymous):

can u help me on like 2 more???

OpenStudy (anonymous):

Do you need more help understanding how we're getting these answers?

OpenStudy (anonymous):

And yeah, I can.

OpenStudy (anonymous):

Jerome walked 2/3 mile in the morning and 11/12 mile in the evening. How many miles did he walk in all?

OpenStudy (anonymous):

2/3 = 8/12 8/12 + 11/12 = 19/12 1 & 7/12 Miles

OpenStudy (anonymous):

Sara had 7/8 pound of turkey. She used 2/3 pound for sandwiches. How many pounds of turkey does Sara have left?

OpenStudy (anonymous):

7/8 = 21 / 24 2/3 = 16/24 She has 5/24 pounds left.

OpenStudy (anonymous):

thanks so much